# A Survey of Constraint Formulations in Safe Reinforcement Learning

### v1

@@ Abstract
Ensuring safety is critical when applying reinforcement learning (RL) to real-world problems. Consequently, safe RL emerges as a fundamental and powerful paradigm for safely optimizing an agent’s policy from experimental data. A popular safe RL approach is based on a constrained criterion, which solves the problem of maximizing expected cumulative reward under safety constraints. Though there has been recently a surge of such attempts to achieve safety in RL, a systematic understanding of the field is difficult due to 1) the diversity of constraint representations and 2) little discussion of their interrelations. To address this knowledge gap, we provide a comprehensive review of representative constraint formulations, along with a curated selection of algorithms specifically designed for each formulation. Furthermore, we elucidate the theoretical underpinnings that reveal the mathematical mutual relations among common problem formulations. We conclude with a discussion of the current state and future directions of safe reinforcement learning research.
@@ 1 Introduction
Reinforcement learning (RL, Sutton and Barto (1998)) is a powerful paradigm for solving sequential decision-making problems through evaluation and improvement, which has made significant progress in many fields such as games (Mnih et al., 2015; Silver et al., 2016), robotics (Levine et al., 2016), and recommendation systems (Afsar et al., 2022). Recently, particular attention has been paid to RL for its use in the fine-tuning of large language models (LLMs) under the name of reinforcement learning from human feedback (RLHF, Ouyang et al. (2022)). RL is a general concept that can be used in many domains and has been spreading to various disciplines as a useful tool.
@@
When applying RL to real problems, safety is an essential requirement (Amodei et al., 2016). Promising areas for safe RL applications include robotics (Pham et al., 2018), autonomous driving (Shalev-Shwartz et al., 2016), and medical care (Jia et al., 2020), and so onL4D. An emerging and crucial use case of safe RL involves refining LLMs using RLHF to align with human preferences. It is critical to prevent harmful outputs (e.g., toxicity, discrimination) while maintaining utility, which is called Safe RLHF (Dai et al., 2023). As such, safe RL has been an active topic in artificial intelligence, which has been studied both theoretically and practically.
@@
Safe RL is a broad concept with different formulations for the different aspects of real-world safety-critical problems. Garcıa and Fernández (2015) provided a survey of safe RL and categorized its optimization criteria into four groups: 1) the constrained criterion (Geibel, 2006), 2) the worst-case criterion (Heger, 1994), 3) the risk-sensitive criterion (Borkar, 2002), and 4) others (e.g., r-squared and value-at-risk). This paper focuses on the constrained criterion of safe RL. It is a growing and powerful group of methods that maximize the expected cumulative reward under safety constraints.
@@
An interesting yet confusing point in the safe RL research based on the constrained criterion is that there are diverse representations of safety constraints yet little discussion regarding the relations among them. Safe RL research in the last decade has focused on developing new algorithms, i.e. how to solve the problem while pursuing performance. New algorithms have continuously been developed under various formulations, thereby making it increasingly challenging to stay abreast of advancements in the field. There are several recent survey papers on safe RL (Brunke et al., 2021; Liu et al., 2021; Gu et al., 2022), but they also focus on methods rather than formulations. As formulation represents the initial phase in comprehending safe RL or implementing algorithms in practical scenarios, it is imperative for the community to comprehensively survey existing literature and lay the groundwork for acquiring a systematic comprehension.
@@
This paper first provides a comprehensive survey focusing on the constraint formulations in safe reinforcement learning and introduces representative algorithms for each formulation. Also, we discuss the relations between each constraint formulation while defining three notions called transformability, generalizability, and conservative approximation. Specifically, we provide theoretical results that there are two problems called identical or more general safe RL (IoMG-SafeRL) problems, into which other common problems can be transformed. The main contribution of this paper is to bridge the problems faced by the reader with appropriate algorithms by organizing existing research with a focus on constraint formulation.
@@ 2 Preliminaries
where $\mathcal{S}$ is a state space $\{s\}$, $\mathcal{A}$ is an action space $\{a\}$, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition, where $\mathcal{P}(s^{\prime}\,|\,s,a)$ is the probability of transition from state $s$ to state $s^{\prime}$ when action $a$ is taken. Note that $\Delta(X)$ denotes the probability simplex over the set $X$. In addition, $H\in\mathbb{Z}_{+}$ is the (fixed) finite length of each episode, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, $\gamma_{r}\in[0,1)$ is the discount factor for the reward, and $\rho\in\Delta(\mathcal{S})$ is the initial state distribution. A key difference from a standard MDP $\mathcal{M}$ lies in the existence of an additional tuple $\mathcal{C}$ to represent safety constraints, which will be referred to as a “constraint tuple” in the rest of this paper. [^1] Though this paper considers finite-horizon discounted CMDPs, key ideas can be extended to infinite and/or undiscounted cases.
@@
An overall sequence for solving safe RL problems based on a constrained criterion is illustrated in Figure 1, which consists of 1) problem formulation and 2) policy optimization. An interesting yet complicated point for understanding safe RL research is that there are several formulations for representing the constrained tuple $\mathcal{C}$ and the feasible policy space $\widehat{\Pi}$ depending on the problem settings or applications that researchers have in their minds. In the next section, we will review seven common safety constraint representations that have been well-studied in existing safe RL research.
@@ 3.6 Almost Surely Instantaneous Safety Constraint with Time-invariant Threshold
Some existing studies formulate safe RL problems via an instantaneous constraint, attempting to ensure safety even during the learning stage while aiming for extremely safety-critical applications such as planetary exploration (Turchetta et al., 2016; Wachi et al., 2018; Wachi and Sui, 2020; Wang et al., 2023). Such studies require the agent to satisfy the following instantaneous safety constraint at every time step.
@@ 3.8 Other Constrained Formulations
Conventional RL algorithms depend on ergodicity assumption; that is, any state $s$ is eventually reachable from any other state $s^{\prime}$ by following a suitable policy. This assumption does not hold in many practical applications since an agent cannot recover on its own after catastrophic actions. Moldovan and Abbeel (2012) removed the ergodicity assumption and proposed an algorithm in which an agent is required to guarantee returnability to the initial safe state.
While this paper has dealt with numerical safety constraints, there have been attempts to represent them using a certain language. For example, Fulton and Platzer (2018) represents the safety constraint via formal languages. This constraint representation is quite useful for leveraging human knowledge in RL, which leads to powerful solutions such as shielding methods (Alshiekh et al., 2018). Finally, Yang et al. (2021) uses natural language to represent safety constraints. This approach is friendly for the general public, which would be a promising method for applying safe RL to real problems.
@@ 4.2 Preliminary Lemmas
4.3 Two IoMG-SafeRL Problems
@@ 5.1 Formulation and Algorithm Selection
As we have discussed in this paper, when we adopt the safe RL paradigm, there are two main steps: 1) problem formulation and 2) policy optimization. Based on the survey results, let us discuss how to solve safe RL problems via constraints.
@@
Another perspective in the algorithm selection should include when an agent needs to satisfy the safety constraint. Existing algorithms can be divided into two classes (see “safety” column in Table 1). The first class (w/o T) tries to achieve the required level of safety after convergence while encouraging safety during training. The algorithms in this class are usually based on Problems 1 or 5 where the safety constraints are represented using the additive safety cost structure. The second class (w/ T) tries to guarantee safety even during training as well as after convergence. The algorithms in this class are typically built upon Problems 6 or 7 where the safety constraints are instantaneous. Which class of algorithm you should choose depends solely on the problem and application to be addressed.
@@
As shown in Table 1, most of the existing safe RL literature considers “online” RL settings where an agent interacts with the environment while learning its policy. A major advantage of safe “online” RL is that policies can be trained from scratch without data collected previously. On the other hand, “offline” RL (Levine et al., 2020) is a framework to train a policy from a fixed amount of pre-collected data, potentially solving the fundamental issues regarding safety or risk. Offline RL is well-suited in the context of safe RL because the agent does not interact with the environment during training; thus, the policy training does not essentially pose any risk. Hence, safe offline RL has been and will be considered a promising and fundamental solution for achieving safety in RL. Although most of the existing safe offline RL literature is based on Problem 1 (e.g., Le et al. (2019), Lee et al. (2021)) as shown in Table 1, such research direction is expected to expand to other problem settings.
@@ 6 Conclusion
This paper provides a survey of safe RL based on constrained criteria from the perspective of formulations. In Section 3, we first present seven common constraint formulations and associated representative algorithms. Also, the curated selection of literature is summarized according to the problem formulation in Table 1. In Section 4, we discuss the theoretical relations among formulations, which will be helpful for readers to obtain a deeper understanding of safe RL. Finally, we describe the current status of safe RL research and potential future directions. With this survey paper, we hope to have presented a systematic understanding and encourage further fundamental and applied research in this area.

### v2

@@ Abstract
Safety is critical when applying reinforcement learning (RL) to real-world problems. As a result, safe RL has emerged as a fundamental and powerful paradigm for optimizing an agent’s policy while incorporating notions of safety. A prevalent safe RL approach is based on a constrained criterion, which seeks to maximize the expected cumulative reward subject to specific safety constraints. Despite recent effort to enhance safety in RL, a systematic understanding of the field remains difficult. This challenge stems from the diversity of constraint representations and little exploration of their interrelations. To bridge this knowledge gap, we present a comprehensive review of representative constraint formulations, along with a curated selection of algorithms designed specifically for each formulation. In addition, we elucidate the theoretical underpinnings that reveal the mathematical mutual relations among common problem formulations. We conclude with a discussion of the current state and future directions of safe reinforcement learning research.
@@ 1 Introduction
Reinforcement learning (RL, (Sutton and Barto, 1998)) is a powerful paradigm for solving sequential decision-making problems through evaluation and improvement, which has made significant progress in many fields such as games (Mnih et al., 2015; Silver et al., 2016; Wurman et al., 2022), robotics (Levine et al., 2016), data center cooling (Li et al., 2019), finance (Hambly et al., 2023), recommendation systems (Afsar et al., 2022), and healthcare (Yu et al., 2021). Recently, particular attention has been paid to RL for its use in the fine-tuning of large language models (LLMs) under the name of reinforcement learning from human feedback (RLHF, Ouyang et al. (2022)). RL is a general concept that can be employed in many domains and has been spreading to various disciplines as a useful paradigm.
@@
When applying RL to real-world problems, safety is an essential requirement (Amodei et al., 2016). Thus, safe RL has been actively studied in recent years so that the benefits of RL are realized while minimizing negative safety issues. Promising areas for safe RL applications include robotics (Pham et al., 2018), autonomous driving (Shalev-Shwartz et al., 2016), healthcare (Jia et al., 2020), and many others. An emerging and crucial use case of safe RL involves refining LLMs using RLHF to align with human preferences. Specifically, it is critical to prevent harmfulness or bias (e.g., toxicity, discrimination) while maintaining the helpfulness of the generated sentences. For example, Safe RLHF (Dai et al., 2023) is a representative approach to balance such a tradeoff. Safe RL is an active area of research in artificial intelligence, with extensive theoretical and practical investigations aimed at developing RL systems that exhibit safe and reliable behavior.
@@
Safe RL is inherently a broad concept with different formulations for the different aspects of real-world safety-critical problems. Garcıa and Fernández (2015) provided an eminent survey of safe RL and categorized its optimization criteria into four groups: 1) constrained criteron (Geibel, 2006), 2) worst-case criteron (Heger, 1994), 3) risk-sensitive criteron (Borkar, 2002), and 4) others (e.g., r-squared, value-at-risk). This paper focuses on the constrained criterion of safe RL as a growing and powerful group of methods to optimize policies under safety constraints.
@@
A noteworthy aspect of safe reinforcement learning research based on constrained optimization criteria is the existence of diverse representations for safety constraints, with little analysis the interrelationships and theoretical connections among these various formulations. Safe RL research in the last decade has focused on developing new algorithms, i.e. how to solve the problem while pursuing performance. New algorithms have continuously been developed under various formulations, thereby making it increasingly challenging to stay abreast of advancements in the field. There are several recent survey papers on safe RL (Kim et al., 2020; Brunke et al., 2021; Liu et al., 2021; Gu et al., 2022; Zhao et al., 2023b), but they also focus on methods rather than formulations. As formulation represents the initial phase in comprehending safe RL or implementing algorithms in practical scenarios, it becomes imperative for the community to comprehensively survey the existing literature and lay the groundwork for acquiring a systematic comprehension.
@@
This paper provides a comprehensive survey focusing on constraint formulations in safe reinforcement learning and introduces representative algorithms for each formulation. Furthermore, we discuss the relationships between various constraint formulations by defining three theoretical notions: transformability, generalizability, and conservative approximation. Specifically, we present theoretical results demonstrating that there exist two problems, termed Identical or More General Safe RL (IoMG-SafeRL) problems, into which other common problems can be either transformed or conservatively approximated. The main contribution of this paper is to bridge the gaps between the safe RL problems with appropriate algorithms by organizing existing research with a focus on constraint formulation.
@@ 2 Preliminaries
where $\mathcal{S}\coloneqq\{s\}$ is a state space, $\mathcal{A}\coloneqq\{a\}$ is an action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition, where $\mathcal{P}(s^{\prime}\,|\,s,a)$ is the probability of transition from state $s$ to state $s^{\prime}$ when action $a$ is taken. Note that $\Delta(X)$ denotes the probability simplex over the set $X$. In addition, $H\in\mathbb{Z}_{+}$ is the (fixed) finite length of each episode, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, $\gamma_{r}\in[0,1)$ is the discount factor for the reward, and $\rho\in\Delta(\mathcal{S})$ is the initial state distribution. A key difference from a standard MDP $\mathcal{M}$ lies in the existence of an additional tuple $\mathcal{C}$ to represent safety constraints, which will be referred to as a “constraint tuple” in the rest of this paper. [^1] Though this paper considers finite-horizon discounted CMDPs, key ideas can be extended to infinite and/or undiscounted cases.
@@
An overall sequence for solving safe RL problems based on a constrained criterion is illustrated in Figure 1, which consists of 1) problem formulation and 2) policy optimization. An interesting yet complicated point for understanding safe RL research is that there are several formulations for representing the constrained tuple $\mathcal{C}$ and the feasible policy space $\widehat{\Pi}$ depending on the problem settings or applications that researchers have in their minds. In the next section, we will review seven common safety constraint representations that have been well-studied in safe RL literature.
@@ 3.6 Almost Surely Instantaneous Safety Constraint with Time-invariant Threshold
Some existing studies formulate safe RL problems via an instantaneous constraint, attempting to ensure safe exploration (Sui et al., 2015) even during the learning stage while aiming for extremely safety-critical RL applications such as planetary exploration and medical treatment (Turchetta et al., 2016; Wachi et al., 2018; Wachi and Sui, 2020; Wang et al., 2023). Such studies require the agent to satisfy the following instantaneous safety constraint at every time step.
@@ 3.8 Other Constrained Formulations
The variance of the return is related to the Sharpe ratio (Sharpe, 1966) – the ratio between the expected profit and its standard deviation. Thus, this formulation is particularly useful in financial appliactions (Meng and Khushi, 2019).
Also, conventional RL algorithms depend on ergodicity assumption; that is, any state $s$ is eventually reachable from any other state $s^{\prime}$ by following a suitable policy. This assumption does not hold in many practical applications since an agent cannot recover on its own after catastrophic actions. Moldovan and Abbeel (2012) removed the ergodicity assumption and proposed an algorithm in which an agent is required to guarantee returnability to the initial safe state.
While this paper has dealt with numerical safety constraints, there have been attempts to represent them using a certain language. For example, Fulton and Platzer (2018) or Hasanbeig et al. (2020) represent the safety constraint via formal languages such as linear temporal logic. This constraint representation is quite useful for leveraging human knowledge in RL, which leads to powerful solutions such as shielding methods (Alshiekh et al., 2018; Li and Bastani, 2020). Finally, Yang et al. (2021) uses natural language to represent safety constraints. Natural language is one of the most powerful mediums yet friendly representations to the general public. Given the recent remarkable progress and utilities of LLMs, this approach would be promising for applying safe RL to real-world AI systems.
@@ 4.2 Preliminary Lemmas
4.3 The Two IoMG-SafeRL Problems
@@ 5.1 Formulation and Algorithm Selection
As mentioned earlier, when we adopt the safe RL paradigm, there are two main steps: 1) problem formulation and 2) policy optimization. Based on the survey results, let us discuss how to solve safe RL problems via constraints.
@@
Another perspective in the algorithm selection should include when an agent needs to satisfy the safety constraint. Existing algorithms can be divided into two classes (see “safety” column in Table 1. The first class (w/o T) tries to achieve the required level of safety after convergence while encouraging safety during training. The algorithms in this class are usually based on Problems 1 or 5 where the safety constraints are represented using the additive safety cost structure. The second class (w/ T) tries to guarantee safety even during training as well as after convergence. The algorithms in this class are typically built upon Problems 6 or 7 where the safety constraints are instantaneous. Which class of algorithm you should choose depends solely on the problem and application to be addressed.
@@
As shown in Table 1, most of the existing safe RL literature considers “online” RL settings where an agent interacts with the environment while learning its policy. A major advantage of safe “online” RL is that policies can be trained from scratch without data collected previously. On the other hand, “offline” RL (Levine et al., 2020) is a framework to train a policy from a fixed amount of pre-collected data, potentially solving the fundamental issues regarding safety or risk. Offline RL is well-suited in the context of safe RL because the agent does not interact with the real environment during training; thus, the policy training does not essentially pose any risk. Hence, safe offline RL is a promising approach to achieve safety in RL if enough data or high-fidelity simulation is available. Although most of the existing safe offline RL literature is based on Problem 1 (e.g., Le et al. (2019), Lee et al. (2021)) as shown in Table 1, such research direction is expected to expand to other problem settings.
@@ 6 Conclusion
This paper provides a comprehensive survey of safe reinforcement learning based on constrained optimization criteria, with a particular focus on problem formulations. We present seven common constraint formulations and their associated representative algorithms in Section 3. Additionally, a curated selection of relevant literature is summarized according to the problem formulation in Table 1. In Section 4, we examine the theoretical relations among these formulations, offering readers a deeper understanding of safe RL. Furthermore, we describe the current status of safe RL research and outline potential future directions. Through this survey paper, we aim to foster a systematic understanding of constraint formulations and encourage further fundamental and applied research in safe reinforcement learning.